Exercise 1 This is the first exercise. Exercise 2 This is the second exercise. Exercise 3 (Category codes…) An exercise on category codes: how to use ’/’ for commands instead of ’\’ ? (For instance, /def/a{macro a}).

Answer of exercise 1 Answer to the first exercise. 2 Answer of exercise 2 Answer to the second exercise.

3 Answer of exercise 3 The “escape” character ’\’ is of category 0. Hence, \catcode`/=0 does the trick.